### v1

# Krylov Complexity in early universe

Abstract
The Lanczos algorithm offers a method for constructing wave functions for both closed and open systems based on their Hamiltonians. Given that the entire early universe is fundamentally an open system, we apply the Lanczos algorithm to investigate Krylov complexity across different phases of the early universe, including inflation, the radiation-dominated period (RD), and the matter-dominated period (MD). Notably, we find that Krylov complexity differs between the closed and open system approaches. To effectively capture the impact of potentials during the RD and MD phases, we analyze various inflationary potentials—including the Higgs potential, the $R^{2}$ inflationary potential, and chaotic inflationary potential—taking into account the violations of slow-roll conditions. This analysis is conducted in terms of conformal time through the preheating process. Our numerical results indicate that the evolution of Krylov complexity and Krylov entropy is remarkably similar across both methods, regardless of the potential under consideration. Additionally, we rigorously construct what is referred to as an open two-mode squeezed state, utilizing the second kind of Meixner polynomials. Based on this construction, we are the first to calculate the evolution equations for $r_{k}$ and $\phi_{k}$ as they relate to the scale factor. Our findings suggest that dissipative effects lead to a rapid decoherence-like behavior. Moreover, our results indicate that inflation behaves as a strongly dissipative system, while both the radiation-dominated and matter-dominated phases exhibit characteristics of weak dissipation. This research provides new insights into exploring the universe from the perspective of quantum information.
arxiv: 1234.56789
1 Introduction
The idea of complexity was initially introduced in the holographic framework Stanford:2014jda . In this framework, it was observed that the conformal field theory (CFT) reaches thermal equilibrium relatively quickly, while the corresponding bulk gravity (wormhole) takes a much longer time to reach the same state. To address this issue, Refs. Hartman:2013qma ; Liu:2013iza introduced the concept of computational complexity within the holographic framework. This concept suggests that while the emergence of spacetime is related to quantum entanglement, the evolution of space occurs through computational complexity.
The concept of complexity has become increasingly essential in high-energy physics even without the holographic principle. First, the so-called computational complexity describes the minimum number of logical steps required for a task Aaronson:2016vto . Furthermore, the method of calculating the complexity was first proposed by Nielson, $\it e.t.c$, dubbed as the geometrical method Nielsen:2005mkt ; Dowling:2006tnk ; Nielsen:2006cea . Another approach is highly related to the ”Fubini study” distance of the geometry of information Chapman:2017rqy . These two geometrical methods are highly dependent on the choice of the geometry of the parametric manifold. For more precisely capturing the information for the quantum state itself, the concept of Krylov complexity was proposed using the Lanczos algorithm viswanath1994recursio , which is explicitly related to the quantum state without needing the parametric manifold. Since this method is free of ambiguities compared with the previous two ones, there will be similarities and differences between them. According to the concept of Krylov complexity, the universal operator growth was proposed Parker:2018yvk . Once the precise quantum state is given, we could analyze its Krylov complexity, Krylov entropy, and Lanczos coefficient based on the Lanczos algorithm viswanath1994recursio . For the connection between the Krylov complexity and the circuit complexity, Ref. Aguilar-Gutierrez:2023nyk has shown that they cannot be compatible with each other. However, Ref. Caputa:2021sib demonstrated that the Krylov complexity was proportional to the distance of the ”Fubuni-Study” metric.
Recently, the Krylov complexity has been broadly applied to condensed matter and high-energy physics. Ref. Muck:2022xfc has systematically analyzed the Krylov complexity in various polynomials dubbed as the orthogonal basis. Due to the importance of the SKY model in high energy physics, the Krylov complexity can also be implemented into this field Rabinovici:2020ryf ; Jian:2020qpp ; He:2022ryk . Meanwhile, the Krylov complexity can be applied to the generalized coherent state Patramanis:2021lkx , Ising and Heisenberg models Cao:2020zls ; Trigueros:2021rwj ; Heveling:2022hth , conformal field theory Dymarsky:2021bjq ; Caputa:2021ori , topological phases of matter Caputa:2022eye and even extend into the quantum field theory He:2024hkw ; He:2024xjp . Ref. Dymarsky:2019elm proved that the chaos can be understood as a delocalization in Krylov complexity. Furthermore, one could observe the exponential growth of Krylov complexity n integrable systems with saddle-dominated scrambling, whose universal features can be verified by Huh:2023jxt . More recent developments of Krylov complexity can be found in Refs. Erdmenger:2023wjg ; Hashimoto:2023swv ; Vasli:2023syq ; Gill:2023umm ; Bhattacharjee:2023uwx ; Adhikari:2022whf ; Loc:2024oen ; Caputa:2024vrn ; Basu:2024tgg ; Sasaki:2024puk ; Caputa:2024xkp ; Bhattacharjee:2022qjw ; Sahu:2024opm ; Kim:2021okd ; Chen:2024imd ; Bhattacharjee:2024yxj ; Sanchez-Garrido:2024pcy ; Chattopadhyay:2024pdj . Especailly, the concept of Krylov complexity can be extended into in various fields Nandy:2024htc .
We aim to investigate the Krylov complexity for the whole early universe, including the inflationary period, radiation-dominated period (RD), and matter-domination period (MD). Some research has examined the computational complexity for inflation Choudhury:2020hil ; Bhargava:2020fhl ; Lehners:2020pem ; Bhattacharyya:2020rpy ; Adhikari:2021ked , where the analysis demonstrated that there is oscillation at the very beginning of inflation, and the evolution was similar in post inflation period. Once considering the quantum effects via various quantum frameworks, the complexity will show the irregular oscillation during inflation providing a new method for distinguishing various inflationary models Li:2023ekd . Quite a few researches are investigating the Krylov complexity in the early universe. Ref. Adhikari:2022oxr is the first one who investigate Krylov complexity in inflation by evaluating the values of non-trivial sound speed, where their method is based on the closed system’s method. However, our universe is an open system. In light of the open system’s method, Ref. Li:2024kfm investigated the Krylov complexity within the framework of modified dispersion relation in inflation, indicating that the inflation behaves as a strong dissipative system and the overall trend of Krylov complexity is increasing. The key ingredient for exploring the Krylov complexity with the open system’s method is constructing the wave function in light of Ref. Bhattacharya:2022gbz , whose wave function wave from Li:2024kfm will serve as our starting point in this work.
In previous studies, Krylov complexity was applied to the early universe using the thermal state. Refs Li:2024ljz ; Li:2024iji examined Krylov complexity by considering the contribution of the potential (quadratic term) and the violation of slow-roll conditions. The potential was approximated to be the quadratic term around the minimum in our previous work Li:2024ljz , which may not capture fully precise information. In the upcoming work, we plan to include more models such as the Higgs potential Higgs:1964pj , the potential of Starobinsky inflation ($R^{2}$ inflation) Starobinsky:1980te , and the potential of chaotic inflation Linde:1983gd . These models can cover most inflationary models, especially those favored by current observations, which indicate that the potential shape is a plateau. It has been suggested by Ref. Galante:2014ifa that many inflationary models can be unified into the so-called $\alpha$ attractors, whose potential is similar to $R^{2}$ inflationary potential. Additionally, Ref. Albrecht:1992kf demonstrated a relation between the Bunch-Davies vacuum and the two-mode squeezed state. We will continue to use the two-mode squeezed state to investigate the Krylov complexity for the entire early universe in light of these two factors.
This paper will be organized as follows. In Sec. 2, we will introduce the Lanczos algorithm. Sec. 3 will give a brief introduction to the early universe, including inflation, RD, and MD, and three different kinds of potentials. Sec. 4 will give the two-mode squeezed state, and the numerical solutions of $r_{k}$ in different periods. Sec. 5 will investigate the Krylov complexity in a method of closed system. Sec. 5.2 will investigate the Krylov entropy in a method of closed system. Sec. 6 will give the Lanczos coefficient and Krylov complexity with the approach of open system, where we are the first to derive the evolution equation of $\phi_{k}$ and $r_{k}$ based on the open two-mode squeezed state. Finally, the conclusion and outlook will be given in Sec. 7.
2 Lanczos algorithm with the approach of closed system
In this section, we will mainly introduce the Lanczos algorithm with the approach of closed system. First, the operator $\mathcal{O}(t)$ in Heisenberg picture needs to be defined as follows,
$\partial_{t}\mathcal{O}(t)=i[H,\mathcal{O}(t)]$ (1)
where H is the Hamilton and its solution is
$\mathcal{O}(t)=e^{iHt}\mathcal{O}e^{-iHt}$ (2)
Next, let’s define the Liouvillian super-operator $\mathcal{L}_{X}$ as $\mathcal{L}_{X}Y=[X,Y]$, where [,][,][ , ] denotes the commutator. We apply this to Eq. (2), the operator can also be expressed as
$\mathcal{O}(t)=e^{i\mathcal{L}t}\mathcal{O}=\sum_{n=0}^{\infty}\frac{(it)^{n}}{n!}\mathcal{L}^{n}\mathcal{O}(0)=\sum_{n=0}^{\infty}\frac{(it)^{n}}{n!}\tilde{\mathcal{O}}_{n}$ (3)
where $\mathcal{L}^{n}\mathcal{O}=\tilde{\mathcal{O}}_{n}=[H,\tilde{O}_{n-1}]$ which can serve as the basis for the Hilbert space.
$\begin{split}\mathcal{O}\equiv|{\tilde{\mathcal{O}}}),\mathcal{L}^{1}\mathcal{O}\equiv|{\tilde{\mathcal{O}}}_{1}),\mathcal{L}^{2}\mathcal{O}\equiv|{\tilde{\mathcal{O}}}_{2}),\mathcal{L}^{3}\mathcal{O}\equiv|{\tilde{\mathcal{O}}}_{3})...\end{split}$ (4)
However, there are not orthogonal bases. In order to construct an orthogonal basis, we use the Lanczos algorithm. The first two operators become
$\mathcal{O}_{0}=|\tilde{\mathcal{O}}_{0})=\mathcal{O},~{}~{}~{}|\mathcal{O}_{1})=b_{1}^{-1}\mathcal{L}|\mathcal{O}_{0})$ (5)
where $b_{1}=\sqrt{(\tilde{\mathcal{O}}_{0}\mathcal{L}|\mathcal{L}\tilde{\mathcal{O}}_{0})}$ desribing the normalized vector. And the next orthogonal basis can be written by
$|\mathcal{O}_{n})=b_{n}^{-1}|A_{n})$ (6)
with
$|A_{n})=\mathcal{L}|\mathcal{O}_{n-1})-b_{n-1}|\mathcal{O}_{n-2}),~{}~{}~{}b_{n}=\sqrt{(A_{n}|A_{n})}$ (7)
where $b_{n}$ represents the Lanczos coefficients. This iterative relation will stop and produce the finite orthogonal Krylov basis if $b_{n}=0$. Subsequently, we can express Eq. (3) as follows,
$\mathcal{O}(t)=e^{i\mathcal{L}t}\mathcal{O}=\sum_{n=0}^{\infty}(i)^{n}\phi_{n}(t)|\mathcal{O}_{n})$ (8)
[…]
$\partial_{t}\phi_{n}(t)=b_{n}\phi_{n-1}-b_{n-1}\phi_{n+1}$ (9)
Next, we can define the Krylov complexity using the wave function $\phi_{n}$ as follows
$K=\sum_{n}n|\phi_{n}|^{2}$ (10)
Additionally, the Lanczos coefficient can be bounded via Parker:2018yvk ,
$b_{n}\leq\alpha n+\eta$ (11)
where $\alpha$ and $\eta$ encodes different information in various models. It should be noticed that $b_{n}=\alpha n+\eta$, meaning that the system is a maximal chaotic dynamical system. For the case of maximal chaos, we could build the relation between the Lyapunov index and $\gamma$ and $\alpha$ as follows,
$\lambda=2\alpha$ (12)
In this work, we will investigate the Krylov complexity for the entire early universe, thus the formula of scale factor $a(\eta)$ should be introduced.
3 Some setup of early universe
In this section, we will first give some basis of scale factor in distinctive periods. Following the notation of our previous work Li:2024ljz , one of the core goal of this work is examining the impact of inflationary potential for the Krylov complexity in RD and MD due to the violation of slow-roll conditions. The Mukhannov-Sasaki variable contains the contribution of potential which cannot show the explicit contribution to the Krylov complexity. That is why we just perturb the inflaton field explicitly in this paper. Thereafter, we will give some very important inflationary potentials, namely, the chaotic inflationary potential, Starobinksy’s inflationary potential ($R^{2}$ inflation) and Higgs potential.
3.1 Inflation, RD and MD
We will follow the notation of Baumann:2009ds to introduce the scale factor. The background metric is the Friedman-Lemaitre-Robertson-Walker metric,
$ds^{2}=a^{2}(\eta)(-d\eta^{2}+d\vec{x}^{2}),$ (13)
where $a(\eta)$ is the scale factor. Here, we will assume a single component universe, which means that the universe will experice the inflation, RD and MD, respectively. We could use equation of state to characterize the universe via
$w_{I}=\frac{P_{I}}{\rho_{I}},$ (14)
where $P_{I}$ and $\rho_{I}$ is the pressure and energy density of the corresponding periods parameterized by $I$. For obtaining the relation between the conformal time and the scale factor, we need the comoving distance,
$\chi_{ph}(\eta)=\int_{t_{i}}^{t}\frac{dt}{a}=\int_{\ln a_{i}}^{\ln a}(aH)^{-1}d\ln a$ (15)
where $\chi_{ph}(\eta)$ is the comoving distance in the event horizon of particle. In light of $w_{I}$, the Hubble radius can be replaced by
$(aH)^{-1}=H_{0}^{-1}a^{\frac{1}{2}(1+3\omega)}$ (16)
where we use $\omega$ to take place of $\omega_{I}$ for simplify. Integrating equation (15) yields the relation between the scale factor and the conformal time
$\eta=\frac{2H_{0}^{-1}}{(1+3\omega)}a^{\frac{1}{2}(1+3\omega)}$ (17)
Consequently, we can derive the relation for various time periods
$\eta=\begin{cases}&-(aH_{0})^{-1},\ \ (\omega=-1),\ \ \rm inflation\\ &aH_{0}^{-1},\ \ (\omega=\frac{1}{3}),\ \ \rm RD\\ &2\sqrt{a}H_{0}^{-1},\ \ (\omega=0),\ \ \rm MD\end{cases}$ (18)
One could see that the various values of $\omega$ describe the distinctive periods.
In the following text, we will discuss the fundamental basis for the different periods in the evolution of the universe. Initially, the universe is created by a singularity, followed by a period of exponential expansion known as inflation. During this phase, the curvature perturbation transitions from the quantum level to the classical level after the horizon exits, leaving an imprint on the Cosmic Microwave Background (CMB). Subsequently, the universe enters the Radiation Dominated (RD) era, characterized by an energy density given by the equation $\rho_{\rm rad}(t)=\rho_{\rm rad_{0}}\frac{a_{0}^{4}}{a(t)^{4}}$, where $\rho_{\rm rad_{0}}$ and $a_{0}$ represent the initial values of the energy density and the scale factor of RD. During the RD era, particles are relativistic, but as the temperature decreases, they become increasingly non-relativistic. This leads the universe into the matter-dominated (MD) epoch, where the energy density is given by $\rho_{\rm mat_{0}}(t)=\rho_{\rm mat_{0}}\frac{a_{0}^{3}}{a(t)^{3}}$. The transition from the RD to the MD epoch occurs when the energy density of radiation becomes equal to the energy density of matter. Once entering MD, there is a process called preheating where the particles will be generated via this process. First, we give the full action for the single field inflation,
[…]
where $V(\phi)$ is the potential of inflation. Following the method of Kofman:1997yn ; Shtanov:1994ce , the equation of motion for background $\phi(t)$ in RD and MD is
$\ddot{\phi}(t)+3H\dot{\phi}(t)+V(\phi),_{\phi}=0,$ (20)
where $\ddot{\phi}(t)=\partial_{t}^{2}\phi(t)$ and the similar work can be found in our previous works Liu:2019xhn ; Liu:2021rgq . Defining with the effective mass $m_{\rm eff}=\frac{d^{2}V}{d\phi^{2}}$ and changing variable as $\tilde{\phi}(t)=a^{3/2}\phi(t)$, Eq. (20) will become
$\ddot{\tilde{\phi}}(t)+m_{\rm eff}^{2}\tilde{\phi}(t)=0,$ (21)
the solution of Eq. (21) is $\phi(t)=a^{-3/2}\cos(m_{\rm eff}t)$ where it is the solution of $\phi(t)$ in preheating period and we have neglecte the term of $\frac{9}{4}H^{2}\tilde{\phi}+\frac{3}{2}\dot{H}\tilde{\phi}$ since it is much smaller compared with potential. It should be noticed that the solution is in terms of physical time $t$. In this paper, we will use $d\eta=\frac{dt}{a}$ to transform into the conformal time for the latter investigations, which is denoted by Eq. (18).
3.2 Various inflationary potentials
There are lots of inflationary models that represent distinctive inflationary potentials. Current observations favor the concave potentials Planck:2018jri , where the potential of $R^{2}$ inflation belongs to this potential. Additionally, the $R^{2}$ inflation agrees with the observations from inflation to large scales. From the theoretical aspect, we could see that many inflationary models can be unified into the so-called $\alpha$ attractors inflationary models Galante:2014ifa , where the other models and $R^{2}$ inflationary model belong to the same kind of inflationary models. Being armed with these two reasons, we will introduce $R^{2}$ inflation. According to the current standard model, the Higgs field is the only scalar field in nature. Therefore, the Higgs potential is our other option Higgs:1964pj . The final choice is the chaotic inflation whose potential is the quadratic term Linde:1983gd . Finally, we could see that our choices could cover the most inflationary models.
Here, we list the formulas of distinctive inflationary potentials as follows,
$\begin{cases}&\ V_{1}=\frac{1}{2}m^{2}\phi^{2},\\ &\ V_{2}=\frac{3}{4}M^{2}M_{p}^{2}(1-e^{-\sqrt{\frac{2}{3}}\frac{\phi}{M_{p}}})^{2},\\ &\ V_{3}=\frac{1}{2}m^{2}\phi^{2}+\frac{1}{4}\lambda\phi^{4},\end{cases}$ (22)
where $M_{p}=(8\pi G)^{-\frac{1}{2}}$ is the definition of the reduced Planck mass. Additionally, $M$ represents the energy scale, and its value can be determined based on the amplitude of the observed power spectrum for primordial perturbations. Being armed with the solution of Eq. (20) and $dt=ad\eta$, we could rewrite $V_{\phi\phi}$ as
$\begin{cases}&\ V_{1,\phi\phi}=m^{2},\\ &\ V_{2,\phi\phi}=M^{2}\Bigl{(}-e^{-\sqrt{\frac{2}{3}}\frac{Aa^{-\frac{3}{2}}\sin(m_{\rm eff}a\eta)}{M_{p}}}+2e^{-2\sqrt{\frac{2}{3}}\frac{Aa^{-\frac{3}{2}}\sin(m_{\rm eff}a\eta)}{M_{p}}}\Bigr{)},\\ &\ V_{3,\phi\phi}=m^{2}+3\lambda A^{2}a^{-3}\sin^{2}(m_{\rm eff}a\eta)\end{cases}$ (23)
where $\phi(t)=Aa^{-\frac{3}{2}}\sin(m_{\text{eff}}t)$. For the effective mass $m_{\text{eff}}$, we have $m_{\text{eff}}^{2}=\frac{d^{2}V(x)}{dx^{2}}$, and it is evident that $m_{\text{eff}}=\sqrt{V_{i,\phi\phi}}$ in our work $(i=1,2,3)$. According to effective mass for the distinctive inflationary models, we need the iterations for $V_{2}$ and $V_{3}$ in terms of $\log_{10}a$ followed by Li:2024kfm . To determine whether the functions $V_{2}$ and $V_{3}$ converge within the defined interval, we performed iterations at multiple sampling points ranging from $\log_{10}a=0$ to $\log_{10}a=6$ with the step $0.05$. Our findings reveal that all points ranging from $\log_{10}a=0$ to $\log_{10}a=6$ are all convergent.
Figure 1: The numeric of $V_{,\phi\phi}$ at first 50 iterations of potential $V_{2}$ and $V_{3}$ in $\log_{10}a=0.2$ and $\log_{10}a=2.8$, in which we set $M_{p}=1$, $m=10^{-6}$; $M^{2}=10^{-12}$ at $V_{2}$, and $\lambda A^{2}=10^{-4}$ at $V_{3}$.
As an illustration, Fig. 1 clearly shows the iterations for potential $V_{2}$ and $V_{3}$, where we have performed $50$ times iterations for two specific values of $\log_{10}a$. Let us mention the range of parameters as shown in Fig. 1, the potential can be constraint by the COBE normalization Planck:2015sxf , whose values are bounded below $10^{-10}$ in Planck units, and the inflationary field is a large field and its maximal value is around $100$. Consequently, the effective mass for various potentials is around $10^{-6}$ and other parameters are set according to this logic.
Figure 2: The numeric of $V_{,\phi\phi}$ in terms of $\log_{10}a$ for potential $V_{2}$, and $V_{3}$ in RD and MD, in which we replace its exact value with the third iteration of $V_{2}$ and $V_{3}$. We set $\theta_{0}=0$, $M_{p}=1$, $m=10^{-6}$; $M^{2}=10^{-12}$ at $V_{2}$, and $\lambda A^{2}=10^{-4}$ at $V_{3}$.
Fig. 2 depicts the evolution of $V_{,\phi\phi}$ with respect to $\log_{10}a$ in RD and MD, using the third iteration of $V_{2}$ and $V_{3}$ via Fig. 1.The evolution of inflation is neglected due to the violation of slow-roll conditions. From Fig. 2, we could see that the effective mass of $R^{2}$ inflation will decrease in RD and then it will grow to the constant ($10^{-12}$) with the oscillations. For the potential of the Higgs field, it is different which shows that there is a peak in RD and after that, it will decrease into some constants. But, it indicates that the effective mass of $V_{2}$ and $V_{2}$ are of almost the same order based on the COBE normalization Planck:2015sxf .
In this section, we provide an introduction to inflation, RD, and MD, which are characterized by the corresponding scale factor. We then present three essential inflationary potentials that encompass most types of inflationary models. Additionally, we provide an overview of preheating, during which the background of the inflaton field oscillates around the minimum. Subsequently, we incorporate the background solution for the inflaton into the numerical calculation of the effective mass for these three distinct potentials.
4 Evolution of two-mode squeezed state
In this section, we will follow the method of Li:2024kfm for investigating the evolution of $r_{k}$ and $\phi_{k}$ in inflation, RD and MD. We have evaluated the quantum circuit complexity in various inflationary frameworks Liu:2021nzx ; Li:2021kfq ; Li:2023ekd .
4.1 Two-mode squeezed state
For obtaining the two-mode squeezed state, we should introduce the unitary operator as follows,
$\mathcal{U}_{k}=\hat{\mathcal{S}}_{k}(\phi_{k},r_{k})\hat{\mathcal{R}}_{k}(\theta_{k})$ (24)
with
$\hat{\mathcal{S}}_{k}=\exp\bigg{[}r_{k}(\eta)\big{(}e^{-2i\phi_{k}(\eta)}\hat{c}_{\vec{k}}\hat{c}_{-\vec{k}}-e^{2i\phi_{k}(\eta)}\hat{c}^{\dagger}_{-\vec{k}}\hat{c}_{\vec{k}}^{\dagger}\big{)}\bigg{]}$ (25)
^$\hat{\mathcal{R}}_{k}(\theta_{k})=\exp\bigg{[}-i\theta_{k}(\eta)\big{(}\hat{c}_{\vec{k}}\hat{c}_{\vec{k}}^{\dagger}+\hat{c}_{-\vec{k}}^{\dagger}\hat{c}_{-\vec{k}}\big{)}\bigg{]}$ (26)
where $\hat{\mathcal{S}}_{k}$ is the squeezed parameter and ^$\hat{\mathcal{R}}_{k}(\theta_{k})$ is the rotation parameter that will be neglected since it only contributes the phase factor of wave function. Then, by applying $\hat{S}_{\vec{k}}(r_{k},\phi_{k})$ to the vacuum state, we can express the wave function of the two-mode squeezed state as follows
$|\psi\rangle=\hat{S}_{\vec{k}}(r_{k},\phi_{k})|0;0\rangle_{\vec{k},-\vec{k}}=\frac{1}{\cosh r_{k}}\sum_{n=0}^{\infty}(-1)^{n}e^{2in\phi_{k}}\tanh^{n}r_{k}|n;n\rangle_{\vec{k},-\vec{k}}.$ (27)
Being armed with this wave function, we could proceed the numeric of $\phi_{k}$ and $r_{k}$. But the evolution of $\phi_{k}$ and $r_{k}$ is based on the total action (19).
4.2 The evolution of $r_{k}(\eta)$ and $\phi_{k}(\eta)$
In this work, we will evaluate the impact of potential for the Krylov complexity, where the Mukhanov-Sasaki variable contains the contribution of potential. Consequently, we cannot observe the explicit influence from potential. Therefore, we will directly perturb the total action into the second order,
[…]
$\pi(\eta,\vec{x})=\frac{\delta L}{\delta f^{\prime}(\eta,\vec{x})}=f^{\prime}-\frac{a^{\prime}}{a}f.$ (29)
According to Eq. (28) and Eq. (29), we could obtain
[…]
Using the Fourier decomposition of operator $\hat{f}(\eta,\vec{x})$ and $\hat{\pi}(\eta,\vec{x})$,
$\hat{f}(\eta,\vec{x})=\int\frac{d^{3}k}{(2\pi)^{3/2}}\sqrt{\frac{1}{2k}}(\hat{c}_{\vec{k}}^{\dagger}f_{\vec{k}}^{*}(\eta)e^{-i\vec{k\cdot}\vec{x}}+\hat{c}_{\vec{k}}f_{\vec{k}}e^{i\vec{k\cdot}\vec{x}}),$ (31)
$\hat{\pi}(\eta,\vec{x})=i\int\frac{d^{3}k}{(2\pi)^{3/2}}\sqrt{\frac{k}{2}}(\hat{c}_{\vec{k}}^{\dagger}u_{\vec{k}}^{*}(\eta)e^{-i\vec{k\cdot}\vec{x}}-\hat{c}_{\vec{k}}u_{\vec{k}}e^{i\vec{k\cdot}\vec{x}}).$ (32)
Substituting equations (31) and (32) into equation (30) to obtain the Hamiltonian operator,
$\hat{H}_{k}=(k+\frac{a^{2}V_{,\phi\phi}}{2k})(\hat{c}_{\vec{k}}\hat{c}_{\vec{k}}^{\dagger}+\hat{c}_{-\vec{k}}^{\dagger}\hat{c}_{\vec{-k}})+(\frac{a^{2}V_{,\phi\phi}}{2k}+i\frac{a^{\prime}}{a})\hat{c}_{-\vec{k}}^{\dagger}\hat{c}_{\vec{k}}^{\dagger}+(\frac{a^{2}V_{,\phi\phi}}{2k}-i\frac{a^{\prime}}{a})\hat{c}_{-\vec{k}}\hat{c}_{\vec{k}}.$ (33)
[…]
$i\frac{d}{d\eta}|\psi\rangle=\hat{H}_{k}|\psi\rangle,$ (34)
we could obtain the evolution of equation of $r_{k}$ and $\phi_{k}$
$\begin{split}&r^{\prime}_{k}=\frac{a^{2}V_{,\phi\phi}}{2k}\sin{2\phi_{k}}-\frac{a^{\prime}}{a}\cos{2\phi_{k}},\\ &\phi_{k}^{{}^{\prime}}=-k-\frac{a^{2}V_{,\phi\phi}}{2k}+(\frac{a^{\prime}}{a}\sin{2\phi_{k}}+\frac{a^{2}V_{,\phi\phi}}{2k}\cos{2\phi_{k}})\coth{2r_{k}}.\end{split}$ (35)
In inflation, Eq. (35) will nicely recover into the case of
$\begin{split}&r_{k}^{{}^{\prime}}=-\frac{a^{\prime}}{a}\cos(2\phi_{k}),\\ &\phi_{k}^{{}^{\prime}}=-k+\frac{a^{\prime}}{a}\sin(2\phi_{k})\coth(2r_{k}).\end{split}$ (36)
It is very difficult to obtain the numeric of Eq. (35), thus we could make a variable transformation, $y=\log_{10}a$, to numerically solve Eq. (35),
$\begin{split}\frac{dr_{k}}{dy}=&\frac{\ln(10)10^{\frac{y}{2}(5+3\omega)}V_{,\phi\phi}}{2kH_{0}}\sin{2\phi_{k}}-\ln(10)\cos{2\phi_{k}}\\ \frac{d\phi_{k}}{dy}=&-\frac{\ln(10)10^{\frac{y}{2}(1+3\omega)}k}{H_{0}}-\frac{\ln(10)10^{\frac{y}{2}(5+3\omega)}V_{,\phi\phi}}{2kH_{0}}+\Bigl{(}\ln(10)\sin{2\phi_{k}}\\ &+\frac{\ln(10)10^{\frac{y}{2}(5+3\omega)}V_{,\phi\phi}}{2kH_{0}}\cos{2\phi_{k}}\Bigr{)}\coth{2r_{k}}\end{split}$ (37)
where we have utilized Eq. (17) and $\frac{a^{\prime}}{a}=H_{0}a^{-\frac{1}{2}(1+3\omega)}$.
Figure 3: The numerical solution of $R_{k}$ in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$. $k=1$ at period inflation, $k=0.01$ at period RD, $k=0.005$ at period MD for simplicity.. The effect of potential is always ignored in slow-roll inflation. But in RD and MD periods, the influence can be seen by using different potential energies as shown in these panels.
Before analyzing the figure of $r_{k}$, it’s important to note that the Krylov complexity depends only on $r_{k}$. Therefore, we only provide the numerical values of $r_{k}$ in Fig. 3 based on Eq. (37). In this equation, we have set $H_{0}=1$ for simplicity and used various values of $k$ in different time periods. The first panel of Fig. 3 illustrates the evolution of $r_{k}$ during inflation. There is an oscillation before the horizon exits, and then $r_{k}$ will grow linearly, consistent with the standard case of Li:2024kfm . As we transition to the radiation-dominated era (RD), we observe that $r_{k}$ will increase to some constant value that oscillates around this value. Interestingly, different potentials produce almost the same trend for $r_{k}$, indicating that the specific potential does not significantly impact the evolution of $r_{k}$. This observation also applies to the matter-dominated era (MD), where the evolution trend is similar to RD. However, we notice that varying momentum significantly affects $r_{k}$, leading to oscillations as the value of $k$ decreases. Although we cannot distinguish between the various potentials using Fig. 3, our numerical data still demonstrates that the evolution of $r_{k}$ is significantly different whether there is a potential or not. It’s worth noting that we cannot provide the values of $r_{k}$ without a potential, as this is not physically meaningful. The potential cannot be ignored due to the violation of slow-roll conditions.
In this section, we have analyzed the evolution of $r_{k}$ for different inflationary potentials within a single-field inflationary framework. Our research indicates that the value of $r_{k}$ increases during inflation. The behavior of $r_{k}$ in radiation-dominated (RD) and matter-dominated (MD) eras is similar and will oscillate around constant values. Our numerical analysis also shows that various inflationary models produce a similar trend for $r_{k}$. Additionally, we have observed that the scale factor and momentum have a significant impact on the evolution of $r_{k}$. All of the details are encoded in Fig. 3.
5 Krylov complexity in closed system
In this section, we will investigate the Krylov complexity in the method of closed system Liu:2021nzx ; Li:2021kfq ; Li:2023ekd , consisting of the Krylov complexity and Krylov entropy.
5.1 Krylov complexity
For obtaining the Krylov complexity, the most essential quantity is the wave function. First, the two-mode squeezed state wave function can be represented by Krylov basis as follows,
$\mathcal{O}(t)=\sum_{n=0}^{\infty}(i)^{n}\phi_{n}(t)|\mathcal{O}_{n})=\frac{1}{\cosh r_{k}}\sum_{n=0}^{\infty}(-1)^{n}e^{2in\phi_{k}}\tanh^{n}r_{k}|n;n\rangle_{\vec{k},-\vec{k}}.$ (38)
Then, we could find the explicit wave function $\phi_{n}$
$\phi_{n}=\frac{1}{\cosh{r_{k}}}(-e^{2i\phi_{k}}\tanh{r_{k}})^{n}.$ (39)
We can define the Krylov compplexity in terms of $\phi_{n}$
$K=\sum_{n}n\left|\phi_{n}\right|^{2}\\ =\cosh^{2}r_{k}\tanh^{2}r_{k}=\sinh^{2}r_{k}.$ (40)
It is explicitly showing that the Krylov complexity is independent of $\phi_{k}$, that is why we only list the figure of $r_{k}$.
Figure 4: The numerical of Krylov complexity in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$. $k=1$ at period inflation, $k=0.01$ at period RD, $k=0.005$ at period MD for simplicity.
Being armed with Eq. (40), we could give its numeric in Fig. 4, indicating the evolution of Krylov complexity in inflation, RD, and MD. Fig. 4 clearly shows that the Krylov complexity is enhanced in inflation which is consistent with the standard case of the modified dispersion within the framework Li:2024kfm . Similar work investigated the Krylov complexity for the exponential expansion for de-Sitter spacetime corresponding to the inflation Adhikari:2022oxr . Ref. Adhikari:2022oxr mentioned that the Krylov complexity equals the average particle number proportional to the volume. Thus we could see that this simple analysis is consistent with the CV conjuncture Stanford:2014jda , in which the complexity is equaled to the volume of wormhole under the framework of holography. However, the method of Adhikari:2022oxr is for the closed system.
In the case of RD and MD, it is observed that the Krylov complexity will increase to certain constant values that fluctuate around this value. Additionally, different patterns of inflationary models tend to produce similar trends of Krylov complexity in different periods. Our analysis aligns with the findings of Barbon:2019wsy , where the Krylov complexity will saturate some constant values even after the scrambling time in a closed system. The scrambling time will be assessed in the context of an open system, particularly with regard to the Lanczos coefficient, Lyapunov index, $\it e.t.c$.
5.2 Krylov Entropy
In a quantum system, entropy is defined as a measure of the disorder of a system. The Krylov entropy (K-entropy) can be used to test the level of disorder of curvature perturbation. Following Barbon:2019wsy , the K-entropy has the following form
$\begin{split}S_{K}&=-\sum_{n=0}^{\infty}\left|\phi_{n}\right|^{2}\ln\left|\phi_{n}\right|^{2}\\ &=\cosh^{2}r_{k}\ln(\cosh^{2}r_{k})-\sinh^{2}r_{k}\ln(\sinh^{2}r_{k})\end{split}$ (41)
where $\phi_{n}$ is the wave function (27). Once given the analytic formula of K-entropy, we could numerically simulate it in various epochs as shown in Fig. 5.
Figure 5: The numerical of Krylov entropy in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$. $k=1$ at period inflation, $k=0.01$ at period RD, $k=0.005$ at period MD for simplicity.
Fig. 5 indicates a complete evolution of K-entropy in inflation, RD and MD, which is similar to Krylov complexity. We could see that the curvature of perturbation will be more and more disorder in inflation. For RD and MD, the disorder will be enhancing and then saturates to some constant values and the total trend is enhancing, where the trend is also highly related the scale factor. From another aspect, the preheating will generate the particles occurred at MD that leads to the growth of K-entropy.
Let’s discuss Section 5. In this section, we provide a comprehensive analysis of Krylov complexity throughout the early universe. Our numerical data clearly indicates that different potentials yield almost identical patterns of Krylov complexity and K-entropy. Specifically, Krylov complexity increases during inflation and then stabilizes at constant values during RD and MD. This same behavior is observed for K-entropy. Additionally, we must examine the Lanczos coefficient (equivalent to the Lyapunov index in maximal chaos), which remains consistent in both the closed and open system methods.
6 Krylov complexity with the approach of open system
The preheating process in MD and RD is a highly non-equilibrium process that creates particles Kofman:1997yn ; Shtanov:1994ce . In addition, Ref. Cheung:2007st demonstrated that the universe’s evolution breaks the time invariance when the temporal direction is reversed, leading to a violation of energy conservation. Due to these reasons, the open system approach is considered to be more realistic compared to the closed system method. We will follow the method of Bhattacharya:2022gbz to investigate the Krylov complexity.
6.1 Lanczos coefficient and dissipation coefficient
Firstly, the general operator in Heisenberg’s picture has the following form,
$\mathcal{O}(\eta)=e^{i\mathcal{L}_{o}\eta}$ (42)
where the Lindbladian $\mathcal{L}_{o}$, also known as the Hamiltonian, acts on the Krylov basis, we can obtain
$\mathcal{L}_{o}|\mathcal{O}_{n})=-ic_{n}|\mathcal{O}_{n})+b_{n+1}|\mathcal{O}_{n+1})+b_{n}|\mathcal{O}_{n-1})$ (43)
where $c_{n}$ encodes the information of the open system and $b_{n}$ is the Lanczos coefficient. Due to this extra term, we could divide the total Hamiltonian (also known as Lindbladian) into two parts which can be represented by
$\mathcal{L}_{o}=\mathcal{H}_{k}=\mathcal{H}_{close}+\mathcal{H}_{open}$ (44)
In our work, $\mathcal{L}_{o}$ is the Hamiltonian Eq. (33). In light of (43), we can easily obtain the open part (generating $c_{n}$) of Hamiltonian,
$\mathcal{H}_{open}=(k+\frac{a^{2}V_{,\phi\phi}}{2k})(\hat{c}_{-\vec{k}}^{\dagger}\hat{c}_{-\vec{k}}+\hat{c}_{\vec{k}}\hat{c}_{\vec{k}}^{\dagger})$ (45)
and the close part (generating $b_{n}$ and $b_{n-1}$) is
$\mathcal{H}_{close}=(\frac{a^{2}V_{,\phi\phi}}{2k}+i\frac{a^{\prime}}{a})\hat{c}_{\vec{k}}^{\dagger}\hat{c}_{-\vec{k}}^{\dagger}+(\frac{a^{2}V_{,\phi\phi}}{2k}-i\frac{a^{\prime}}{a})\hat{c}_{\vec{k}}\hat{c}_{-\vec{k}}$ (46)
where $|\mathcal{O}_{n})=\Bigl{|}n_{\vec{k}};n_{-\vec{k}}\Bigr{\rangle}$. By substituting Eq. (45) and Eq. (46) into Eq. (43), we can obtain the equations for $c_{n}$ and $b_{n}$ as
$c_{n}=i(2n+1)(k+\frac{a^{2}V_{,\phi\phi}}{2k}).$ (47)
$b_{n}=n\sqrt{[(\frac{a^{2}V_{,\phi\phi}}{2k})^{2}+(\frac{a^{\prime}}{a})^{2}]}.$ (48)
Here, we will discuss more about $b_{n}$ and $c_{n}$. Ref. Bhattacharjee:2022lzy has introduced the $a_{n}$ corresponding to $c_{n}$ in our work and also stated that $c_{n}$ is highly related to the dissipative coefficient. Meanwhile, we could see that $b_{n}$ is the same for the closed system and open system since it is precisely determined by the Hamiltonian.
In Ref. Parker:2018yvk , it has clearly indicated that the dynamical system is an infinite, non-integrable, many-body system as $b_{n}=\alpha n+\gamma$, where $\alpha$ and $\gamma$ are two parameters determined by various models. In our case, we could see that $b_{n}\propto n$, thus the whole early universe is an ideal chaotic system.
6.1.1 $b_{n}$, $c_{n}$ and $\mu_{2}$
[…]
$\partial_{\eta}\phi+2b_{n}\partial_{n}\phi+\tilde{c}_{n}\phi=0,$ (49)
where we have defined $\tilde{c}_{n}=-ic_{n}$, meanwhile we have used the $b_{n+1}\approx b_{n}$ and $\phi_{n+1}-\phi_{n-1}\approx 2\partial_{n}\phi$ in the continous limit. $b_{n}$ and $c_{n}$ are defined in Eq. (47) and Eq. (48), respectively. The generic formulas of $b_{n}$ and $c_{n}$ are
$b_{n}^{2}=|1-\mu_{1}^{2}|n(n-1+\beta),\ \ c_{n}=i\mu_{2}(2n+\beta)$ (50)
where $\mu_{1}$, $\mu_{2}$ and $\beta$ are determined by various models and $\mu_{2}$ plays a role of dissipative coefficient as shown in Ref. Li:2024kfm . In our work, one can explicitly find $\beta=1$ and combine Eqs. (47), (48), and (50), we find that the parameters $\mu_{1}$ and $\mu_{2}$ have the following forms,
$|1-\mu_{1}^{2}|=(\frac{a^{2}V_{,\phi\phi}}{2k})^{2}+(\frac{a^{\prime}}{a})^{2},\mu_{2}=k+\frac{a^{2}V_{,\phi\phi}}{2k}.$ (51)
It should be noticed that $\mu_{2}$ is mainly determined by $k$ in inflation due to the slow-roll conditions. Once entering RD and MD (taking the contribution of the effective mass into account), we could see that it will be smaller than unity since the effective mass of these three types potentials is much more tiny compared with momentum $k$. First, we discuss the Lanczos coefficient $b_{n}$. Fig. 6 clearly indicates the three distinctive potentials of $b_{n}$ in inflation, RD and MD, in which the parameters of $k$, $H_{0}$ are set to be the same with the previous plots. As Bhattacharya:2022gbz confirmed, the Lyapunov index is equavillent to $b_{n}$ whose relation is $\lambda=2\alpha$ with $b_{n}=\alpha n+\delta$ ($\delta=0$ in our case), where it is explicitly showing that $\lambda=2\alpha$, namely $\lambda=b_{n}$ with $n=2$ which will be shown in Fig. 6. The physical meaning of $\lambda$ represents the chaotic feature of dynamical system. Consequently, $b_{n}$ in our case also manifest the chaotic feature for the early universe.
Figure 6: The plot of $b_{n}=\lambda$ with $n=2$.The numerical of $b_{n}$ in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$. $k=1$ at period inflation, $k=0.01$ at period RD, $k=0.005$ at period MD for simplicity.
Figure 6 illustrates the evolution trend $b_{n}$ in inflation, radiation domination (RD), and matter domination (MD). Based on the discussion regarding $b_{n}$, chaos will be heightened due to the exponential expansion of the background space. As the energy decreases, $b_{n}$ will also decline since the temperature decreases. During MD, particles are generated through a preheating process known as parametric resonance, which leads to an increase in $b_{n}$. Finally, we need to analyze the dissipative coefficient $\mu_{2}$ that is related to $c_{n}$, where we could confirm the previous results via Li:2024kfm . Fig. 7 clearly shows the evolution in RD and MD. The $\mu_{2}$ in inflation is not shown in Fig. 7 since $\mu_{2}=k=1$ in light of the values of $k$ for inflation. Ref. Li:2024kfm claimed that inflation is a strong dissipative system as $\mu_{2}\geq 1$, RD and MD are weak dissipative systems as $\mu_{2}\ll 1$, which is explicitly confirmed by Fig. 7.
To summarize the chaotic features of the early universe: First, chaos is enhanced by the exponential expansion. Then, as the energy scale diminishes, chaos decreases. Finally, the generation of particles during MD through the preheating process increases chaos once again.
Figure 7: The numerical of dissipative coefficient $\mu_{2}$ in terms of $\log_{10}a$ for RD and MD, where we set $H_{0}=1$, $k=0.01$ at period RD, $k=0.005$ at period MD for simplicity. The value of $mu_{2}$ indicates the RD and MD are weak dissipative systems.
6.2 General discussion of $\phi_{n}$
Before the Krylov complexity, we discuss the general properties of $\phi_{n}$ since the definition of Krylov complexity is explicitly represented by $C_{K}=\sum_{n=1}^{+\infty}n|\phi_{n}|^{2}$. We will follow the general discussion of $\phi_{n}$ via Refs. Li:2024kfm ; Bhattacharjee:2022lzy . Eq. (49) can be rewritten by
$\partial_{\eta}\phi+n(\chi\mu\phi+2\alpha\partial_{n}\phi)=0,$ (52)
where we have defined $\tilde{c}_{n}=n\chi\mu$ and $b_{n}=n\alpha$. As $\eta$ approaches infinity, the solution of Eq. (52) has the form $\phi(n)\propto e^{-n\xi}$ with $\xi=\frac{2\alpha}{\mu\chi}$, where this form will appear in RD or MD due to their weak dissipative nature. In the weak dissipative region, the corresponding Krylov complexity behaves as follows: it increases exponentially at the onset, with the average position given by $C_{k}\propto e^{2\alpha\eta}$ before reaching $\eta_{*}$. After this point, the Krylov complexity saturates to a constant value $\xi$ when $\eta>\eta_{*}$, where $\eta_{*}\sim\frac{1}{2\alpha}\ln{\frac{2\alpha}{\chi\mu}}$. It is straightforward to demonstrate that $\exp(2a\eta_{*})=\xi$. After some simple algebra, we could obtain the formula $\xi$ in our case as follows,
$\xi=\frac{2\sqrt{(\frac{a^{2}V_{,\phi\phi}}{2k})^{2}+(\frac{a^{\prime}}{a})^{2}}}{(2+\frac{1}{n})(k+\frac{a^{2}V_{,\phi\phi}}{2k})}$ (53)
From this formula, we can observe that $\xi$ is not strictly constant; it varies with respect to $a$ and $k$. Consequently, a significant change in the background will have a substantial impact on Krylov complexity. In this section, we examined the general properties of Krylov complexity without providing the precise formula for $\phi_{n}$. To fully capture the information concerning Krylov complexity within the framework of the open system method, it is essential to have the exact wave function for $\phi_{n}$.
6.3 Wave function within the method of open system
Although we constructed the wave function for an open system in reference Li:2024kfm , based on the framework established in Bhattacharjee:2022lzy , our approach is too simplistic to capture the complete information of the open system. Notably, our wave function, as shown in Li:2024kfm , can effectively reduce to the case of a closed system, establishing a correspondence between $\eta$ and $r_{k}$. To clearly illustrate this issue, we will first present our constructed wave function as follows:
$\phi_{n}=\frac{\rm sech\eta}{1+\mu_{2}\tanh\eta}|1-\mu_{1}^{2}|^{\frac{n}{2}}(\frac{\tanh\eta}{1+\mu_{2}\tanh\eta})^{n}.$ (54)
To perform the Taylor expansion around $\mu_{2}$ to the first order, we obtain the expression
$\phi_{n}\approx|1-\mu_{1}^{2}|^{n/2}\,\text{sech}(\eta)\,\tanh^{n}(\eta)+\mathcal{O}(\mu_{2}^{n}).$
Next, we can neglect the term involving $\mu_{1}^{2}$, which enables us to find that this approximation matches the wave function given in equation (39) for $\phi_{k}=0$. Additionally, we observe the correspondence $\eta\sim r_{k}$. In principle, the wave function $\phi_{n}$ should incorporate the information of both $r_{k}$ and $\phi_{k}$ to accurately reproduce the wave function represented in (39) as the leading order of the open system’s method. To achieve this, we will follow the guidelines provided in Appendix D of Parker:2018yvk .
For deriving the two-mode squeezed state in open system, we first introduce the concept of orthogonal polynomial sequence (OPS) $\{P_{n}(x)\}_{n\geq 0}$, it satisfied with the following recurrence relation,
$P_{n+1}(x)=(x-\tilde{c}_{n})P_{n}(x)-b_{n}^{2}P_{n-1}(x),\ \ \ \ \mbox{for}\ \ n\geq 1$ (55)
where $x$ represents the Hamiltonian, $\tilde{c}_{n}$ is associated with the dissipative coefficient, and $b_{n}$ refers to the Lanczos coefficient. The sequence begins with the initial conditions $P_{0}(x)=1$ and $P_{1}(x)=x-c_{0}$. For an open system, we have $P_{n}(x;\alpha,\eta)=\det(x-\mathcal{L}_{n})$, where $\mathcal{L}_{n}$ is the Hamiltonian for the $n$-th quantum state. This determinant encodes information about $b_{n}$ as represented by the corresponding matrix. Based on the definition of $L$, we can express it in terms of $x^{n}$ or $e_{n}$, which will be used to derive the wave function using the open system method.
To obtain the similar recurrence relation as (43), we introduce the the natural orthonormal basis $\{e_{n}\}$, and meanwhile $P_{n}$ can be represented in terms of $e_{n}$
$P_{n}(x)=\Bigl{(}\prod_{k=1}^{n}b_{k}\Bigr{)}e_{n},\ \ \ \ \mbox{and}\ \ \ x^{n}=\mathcal{L}^{n}e_{0},$ (56)
where $x$ is the parameter of $P_{n}$, $\mathcal{L}$ represents our Hamiltonian expressed in terms of creation and annihilation operators, and $|e_{n}\rangle\equiv|\mathcal{O}_{n}\rangle=\Bigl{|}n_{\vec{k}};n_{-\vec{k}}\Bigr{\rangle}$. By combining Equation (55) with Equation (56), we can derive the following recurrence relation.
$b_{n+1}e_{n+1}+b_{n}e_{n-1}=(x-\tilde{c}_{n})e_{n}.$ (57)
Here, $x$ functions as the Hamiltonian. Given the recurrence of $P_{n}$, it aligns with the second kind of Meixner polynomials, denoted as $M_{n}(x;\delta,\eta_{1})$. Its generating function is
$\sum_{n=0}^{\infty}M_{n}(x;\delta,\eta_{1})\frac{t^{n}}{n!}=((1+\delta t^{2})+t^{2})^{-\eta/2}\exp(x\arctan(\frac{t}{1+\delta t}))$ (58)
The inverse polynomial of $\{M_{n}(x;\delta,\eta_{1})\}_{n\geq 0}$ is $\{Q_{n}(x;\delta,\eta_{1})\}_{n\geq 0}$, which is defined as
$Q_{n}(x)=\sum_{k=0}^{n}\mu_{n,k}x^{k},\ \ \ \ \mbox{and}\ \ \ x^{n}=\sum_{k=0}^{n}\mu_{n,k}P_{k}(x).$ (59)
where $\mu_{n,k}$ refers to a triangular linear transform with matrix elements that relate $x$ to $P_{n}$. This formula demonstrates that the polynomials $P_{n}$ can be expressed in terms of $x$. All three orthogonal polynomial sequences (OPS) mentioned are examples of Sheffer orthogonal polynomials, which are defined as a set of polynomials $\{P_{n}(x)\}_{n\geq 0}$ characterized by a generating function,
$\sum_{n=0}^{\infty}P_{n}(x)\frac{t^{n}}{n!}=f(t)\exp(xg(t)).$ (60)
And the inverse polynomials are give by
$\sum_{n=0}^{\infty}Q_{n}(x)\frac{t^{n}}{n!}=\frac{1}{f(g^{\langle-1\rangle}(t))}\exp(xg^{\langle-1\rangle}(t)),$ (61)
where $g^{\langle-1\rangle}(t)$ stands for the compositional inverse of $g(t)$, which have the following relationship
$g\bigl{(}g^{\langle-1\rangle}(t)\bigr{)}=t.$ (62)
Back to the Meixner polynomials of the second kind, we find that
$f(t)=((1+\delta t)^{2}+t^{2})^{-\eta/2},\ \ \ \ \mbox{and}\ \ \ g(t)=\arctan(\frac{t}{1+\delta t}).$ (63)
Our target function is $Q_{n}(x)$, and the rationale will be provided in the final derivations. To obtain $Q_{n}(x)$, we need to know the formula for $g^{\langle-1\rangle}$. The compositional inverse polynomial of $g(t)$ is
$t=g\bigl{(}g^{\langle-1\rangle}(t)\bigr{)}=\arctan(\frac{g^{\langle-1\rangle}(t)}{1+\delta g^{\langle-1\rangle}(t)}),$ (64)
Next,
$\frac{g^{\langle-1\rangle}(t)}{1+\delta g^{\langle-1\rangle}(t)}=\tan(t)$ (65)
And after a few simple calculations,
$g^{\langle-1\rangle}(t)=\frac{\tan(t)}{1-\delta\tan(t)},$ (66)
where $\delta$ is a parameter determined by various models. Combine Eqs. (66) and (61), we obtain the inverse polynomials of $\{M_{n}(x;\delta,\eta_{1})\}_{n\geq 0}$ as following,
$\sum_{n=0}^{\infty}Q_{n}(x)\frac{t^{n}}{n!}=\frac{\sec(t)^{\eta_{1}}}{(1-\delta\tan(t))^{\eta_{1}}}\exp(x\frac{\tan(t)}{1-\delta\tan(t)}).$ (67)
After the changing the variable as $t=i\eta$ and $\delta=i\mu_{2}$, where $\eta$ (conformal time) and $\mu_{2}$ (dissipation coefficient) are our defined parameters. Making use of $\tan i\eta=i\tanh\eta$ and $\cos i\eta=\cosh\eta$, the generating function of the inverse polynomials become
$\sum_{n=0}^{\infty}Q_{n}(x)\frac{(i\eta)^{n}}{n!}=\frac{\rm sech(\eta)^{\eta_{1}}}{(1+\mu_{2}\tanh(\eta))^{\eta_{1}}}\exp(x\frac{i\tanh(\eta)}{1+\mu_{2}\tanh(\eta)}).$ (68)
Based on Eqs. (56) and (59), we could obtain
$(e_{d}|\mathcal{L}^{n}|e_{0})=\mu_{n,d}\prod_{k=1}^{n}b_{k}.$ (69)
Furthermore, we could have
$(e_{d}|e^{i\mathcal{L}\eta}|e_{0})=\prod_{k=1}^{d}b_{k}\sum_{n=0}^{\infty}\frac{(i\eta)^{n}}{n!}\mu_{n,d}.$ (70)
In light of (59), we could obtain
$\sum_{n=0}^{\infty}\frac{(i\eta)^{n}}{n!}\mu_{n,d}=\frac{\rm sech\eta}{1+\mu_{2}\tanh\eta}\frac{1}{d!}\frac{(i\tanh\eta)^{d}}{(1+\mu_{2}\tanh\eta)^{d}},$ (71)
Then we utilize $b_{n}=\alpha n=\sqrt{|1-\mu_{1}^{2}|}n$ in our case, we naturally get the following formula,
$(e_{n}|e^{i\mathcal{L}\eta}|e_{0})=\frac{\rm sech\eta}{1+\mu_{2}\tanh\eta}|1-\mu_{1}^{2}|^{\frac{n}{2}}\frac{(i\tanh\eta)^{n}}{(1+\mu_{2}\tanh\eta)^{n}}$ (72)
thus, we obtain the wave function
$|\mathcal{O}(t))=e^{i\mathcal{L}\eta}|e_{0})=\frac{\rm sech\eta}{1+\mu_{2}\tanh\eta}\sum_{n=0}^{\infty}|1-\mu_{2}^{2}|^{\frac{n}{2}}\frac{(-1\exp(2i\phi_{k}(\eta))\tanh\eta)^{n}}{(1+\mu_{2}\tanh\eta)^{n}}|e_{n})$
where
$|\mathcal{O}_{n})=|e_{n})$
In our calculations, we multiply a phase factor $ie^{2i\phi_{k}(\eta)}$ which does not alter the wave function. Finally, we can set $r_{k}=\eta$, allowing us to denote the two-mode squeezed state within the open system as,
$\phi_{n}(\eta)=(-1)^{n}e^{2in\phi_{k}}\frac{\rm sechr_{k}}{1+\mu_{2}\tanh r_{k}}|1-\mu_{1}^{2}|^{\frac{n}{2}}\frac{(\tanh r_{k})^{n}}{(1+\mu_{2}\tanh r_{k})^{n}}.$ (73)
This wave function is our central result in this paper whose leading order of $\phi_{n}$ in terms of $\mu_{2}$ is consistent with the wave function described in (39). The idea of changing the variable is borrowed from the construction of coherent states, where a key component is the generation of the displacement operator, defined as $D(\eta)=e^{iHt}$, with $H$ being the corresponding Hamiltonian, where one can change $t$ in terms of $r_{k}$ and $\phi_{k}$. In Ref. Caputa:2021sib , it is stated that the generalized coherent state is derived by constructing a generalized displacement operator according to their group representation theory.
6.4 The power of wave function (73)
We discuss the two-mode squeezed state in an open system, as described in equation (73). Unlike the two-mode squeezed state derived from closed system methods, our wave function (73) is more realistic, as it considers a vast open environment. Additionally, although our derivation pertains to the early universe, this wave function can also be applied to fields such as quantum optics, quantum information, and condensed matter physics, where information is encoded in the parameters $\mu_{1}$ and $\mu_{2}$ determined by the specific Hamiltonian. Finally, we observe that our calculations hold true by requiring that the recurrence relation (55) which is equivalent to the Lanczos algorithm. This connection highlights the power of the Lanczos algorithm, which provides a natural method for constructing the wave function according to their Hamiltonian, even for open systems.
6.5 The evolution of $r_{k}(\eta)$ and $\phi_{k}(\eta)$ in open system’s method
According to the standard procedure, we need to derive the evolution equations for $r_{k}$ and $\phi_{k}$ in relation to conformal time $\eta$. The detailed calculations for this derivation can be found in Appendix A. The evolution can be summarized as follows:
$\begin{split}r_{k}^{\prime}=&\ \frac{|1-\mu_{1}^{2}|^{\frac{1}{2}}\sinh 2r_{k}\bigl{(}\frac{a^{2}V_{,\phi\phi}}{2k}\sin 2\phi_{k}-\frac{a^{\prime}}{a}\cos 2\phi_{k}\bigr{)}-\sinh 2r_{k}\mu_{2}^{\prime}}{(\sinh 2r_{k}+2\mu_{2}\cosh^{2}r_{k})}\\ \phi_{k}^{\prime}=&\ -\bigl{(}k+\frac{a^{2}V_{,\phi\phi}}{2k}\bigr{)}+\frac{1}{2}\bigl{(}\frac{a^{2}V_{,\phi\phi}}{2k}\cos 2\phi_{k}+\frac{a^{\prime}}{a}\sin 2\phi_{k}\bigr{)}\\ &\ \Bigl{[}|1-\mu_{1}^{2}|^{\frac{1}{2}}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}+|1-\mu_{1}^{2}|^{-\frac{1}{2}}(\tanh^{-1}r_{k}+\mu_{2})\Bigr{]},\end{split}$ (74)
where prime denotes the varying with respect to the conformal time $\eta$. The same as the closed system, we change variable as $y=\log_{10}a$, the evolution equation of $r_{k}$ and $\phi_{k}$ will become as,
$\begin{split}\frac{H_{0}}{\ln 10}10^{-\frac{y}{2}(1+3\omega)}\frac{dr_{k}}{dy}=&\ \frac{|1-\mu_{1}^{2}|^{\frac{1}{2}}\sinh 2r_{k}\bigl{(}\frac{10^{2y}V_{,\phi\phi}}{2k}\sin 2\phi_{k}-H_{0}10^{-\frac{y}{2}(1+3\omega)}\cos 2\phi_{k}\bigr{)}-\sinh 2r_{k}\mu_{2}^{\prime}}{(\sinh 2r_{k}+2\mu_{2}\cosh^{2}r_{k})}\\ \frac{H_{0}}{\ln 10}10^{-\frac{y}{2}(1+3\omega)}\frac{d\phi_{k}}{dy}=&\ -\bigl{(}k+\frac{10^{2y}V_{,\phi\phi}}{2k}\bigr{)}+\frac{1}{2}\bigl{(}\frac{10^{2y}V_{,\phi\phi}}{2k}\cos 2\phi_{k}+H_{0}10^{-\frac{y}{2}(1+3\omega)}\sin 2\phi_{k}\bigr{)}\\ &\ \Bigl{[}|1-\mu_{1}^{2}|^{\frac{1}{2}}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}+|1-\mu_{1}^{2}|^{-\frac{1}{2}}(\tanh^{-1}r_{k}+\mu_{2})\Bigr{]}\end{split}$ (75)
where
$\begin{split}&\ \mu_{2}=k+\frac{10^{2y}V_{,\phi\phi}}{2k},\ \ \ \ \mu_{2}^{\prime}=H_{0}10^{-\frac{y}{2}(1+3\omega)}\frac{10^{2y}V_{,\phi\phi}}{k}+\frac{10^{2y}V_{,\phi\phi}^{\prime}}{2k}\\ &\ |1-\mu_{1}^{2}|=(\frac{10^{2y}V_{,\phi\phi}}{2k})^{2}+(H_{0}10^{-\frac{y}{2}(1+3\omega)})^{2}.\end{split}$
During the calculations, we accounted for the background of $\phi$ using conformal time $\eta$, as demonstrated by the various potentials in Eq. (23). Consequently, $\mu_{2}$ and $\mu_{1}$ also vary with respect to conformal time, since they include contributions from the potential and the scale factor $a$.
Figure 8: The numerical of $r_{k}$ in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$. $k=1$ at period inflation, $k=0.01$ at period RD, $k=0.005$ at period MD for simplicity.
Fig. 8 clearly illustrates the evolution of $r_{k}$ across various epochs, including inflation, radiation domination (RD), and matter domination (MD). It is evident that the different potentials produce similar trends for $r_{k}$. Unlike in Fig. 3, there is no oscillation of $r_{k}$ once dissipation effects are considered. Additionally, the trend of $r_{k}$ varies, particularly during the MD phase, where the evolution consistently shows an increase. These differences will result in distinctive evolutions of Krylov complexity and Krylov entropy.
6.6 Krylov complexity and Krylov entropy with the approach of open system
In this section, we will continue use the wave function (73) to explore the Krylov complexity and Krylov entropy. All of these calculation are exactly the same with our previous study Li:2024kfm , the only differences is that we need to use $r_{k}$ to take place of $\eta$ in the formulas for Krylov complexity and Krylov entropy whose formulas can be expressed as following:
$K=\frac{\rm sech^{2}r_{k}|1-\mu_{1}^{2}|\tanh^{2}r_{k}}{[1+2\mu_{2}\tanh r_{k}+(\mu_{2}^{2}-|1-\mu_{1}^{2}|)\tanh^{2}r_{k}]^{2}}$ (76)
and
$\begin{split}S_{k}=&\ \frac{\rm sech^{2}r_{k}}{A^{2}}\bigl{[}(1+\mu_{2}\tanh r_{k})^{2}\ln(1+\mu_{2}\tanh r_{k})^{2}-\tanh^{2}r_{k}|1-\mu_{1}^{2}|\ln\tanh^{2}r_{k}\\ &\ -[(1+\mu_{2}\tanh r_{k})^{2}-|1-\mu_{1}^{2}|\tanh^{2}r_{k}]\ln\rm sech^{2}r_{k}-\tanh^{2}r_{k}|1-\mu_{1}^{2}|\ln|1-\mu_{1}^{2}|\bigr{]}\end{split}$ (77)
where we have defined $A=1+2\mu_{2}\tanh r_{k}+(\mu_{2}^{2}-|1-\mu_{1}^{2}|)\tanh^{2}r_{k}$ for simplicity. It straightforwardly demonstrates the leading order of Krylov complexity (76) and Krylov entropy (77) will nicely recover into the case of Eq. (40) and Eq. (41) under the weak dissipative approximation, namely, $\mu_{2}\ll 1$.
Fig. 9 illustrates the evolution of Krylov complexity during inflation, radiation domination (RD), and matter domination (MD). During inflation, we observe that Krylov complexity increases exponentially, which is consistent with the closed system’s method. However, the trends in RD and MD differ from those shown in Fig. 4, where the complexity tends to stabilize around certain constant values, oscillating around these values in both RD and MD. In the case of an open system, it is evident that Krylov complexity decreases during RD. For MD, Krylov complexity initially peaks as the scale factor $a$ increases and then declines to stabilize at constant values. Additionally, it is noteworthy that the order of Krylov complexity in an open system is significantly lower than in a closed system, even in RD and MD. Our numerical analysis shows that the term $\rm sech(r_{k})$ in Eq. (76) leads to this outcome, regardless of whether the system is weakly dissipative or not. We have demonstrated that both RD and MD can be approximated as weakly dissipative systems under the condition $\mu<1$. Recalling the meaning of Krylov complexity, it describes the operator growth in a many-body system and our early universe is a perfect chaotic system due to $b_{n}=\alpha n$. Our numeric reveals that the operator growth is much smaller compared with the case of a closed system, which is consistent with reality since the dissipation will lead to faster decreasing of operator growth compared with that without the dissipation effects.
Once one understands the role of dissipation effects, it becomes evident that the order of Krylov entropy in an open system is lower compared to that in a closed system, as illustrated in Figs 10 and 5. The evolution trend is particularly distinctive for RD and MD, as shown in Figure 10, where the Krylov entropy initially peaks before stabilizing at constant values. In the case of MD, the Krylov entropy initially holds steady at certain values before decreasing to some constant level. Overall, the background conditions and dissipative effects have a significant impact on both Krylov complexity and Krylov entropy.
In this section, we utilized the wave function presented in Eq. (73) to investigate Krylov complexity and Krylov entropy. By comparing our findings to those from a closed system, we observed that the dissipative effects not only influence the overall evolution of the system but also reduce the order of Krylov complexity and Krylov entropy. Since Krylov complexity reflects the growth of operators—a fundamental quantum property of a dynamical system—we can conclude that, when the growth of the operator is diminished in the open system compared to the closed system, the dissipative effects lead to a more rapid decoherence-like behavior. Meanwhile, we also could see that the various potentials will almost produce the same trend of Krylov complexity and Krylov entropy in different epochs.
Figure 9: The numerical of Krylov complexity in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$, $k=0.01$ at period inflation and RD, $k=0.005$ at period MD for simplicity.
Figure 10: The numerical of Krylov entropy in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$, $k=0.01$ at period inflation and RD, $k=0.005$ at period MD for simplicity.
7 Summary and outlook
The definition of Krylov complexity is clear and unambiguous, especially when compared to circuit complexity. Its uniqueness eliminates the need for introducing a parametric manifold to illustrate the concept of complexity. Additionally, Krylov complexity characterizes the growth of operations within a dynamical system, and the order of Krylov complexity can indicate the quantum level of that system. Krylov complexity is rooted in the Lanczos algorithm, a powerful method for assessing the nature of dynamical systems, which can be chaotic, integrable, or free. The key criteria for this assessment is the Lanczos coefficient, whose formula indicates the characteristics of a given dynamical system. For example, the relation $b=\alpha n$ suggests that the system is chaotic. In our paper, we highlight that the Lanczos coefficient effectively reveals that our early universe was a many-body, infinite, and chaotic system. From a cosmological perspective, the universe is considered an entirely open system. Notably, the Lanczos algorithm offers a natural way to construct the wave function of an open system using the second kind of Meixner polynomials. Given these insights, it is particularly important to explore Krylov complexity throughout the early universe, including inflation, RD and MD.
In this paper, we thoroughly examine Krylov complexity in both open and closed system methods. Our primary result is the derivation of the two-mode squeezed state using the open system approach, along with the calculation of its evolution equations for $\phi_{k}(\eta)$ and $r_{k}(\eta)$. Below, we present our key findings.
$(a)$. In Sec. 2, we introduce the Lanczos algorithm and the associated Lanczos coefficients, which apply equally to both open and closed systems. As noted by Parker:2018yvk , the Liouvillian superoperator acts as the Hamiltonian when it can be expressed in terms of creation and annihilation operators. Our exploration of Krylov complexity encompasses the entire universe, including inflation, RD, MD, and the violations of slow-roll conditions, which necessitates considering contributions from potential energy. Sec. 3 discusses the most common potentials, such as the Higgs potential, Starobinsky’s inflationary potential, and chaotic inflationary potential. These models encompass a wide range of inflationary scenarios due to the unifying characteristics of $\alpha$ attractors. To accurately capture the information pertaining to these potentials, we examine the dynamics of the field $\phi$ in relation to the scale factor during the preheating process, as represented in Eq. (21). From these calculations, we determine the effective mass for various inflationary potentials, as illustrated in Fig. 2. Our findings indicate that the potential approaches certain constant values during RD and MD.
$(b)$. In light of the scale factor and various potentials, we examine Krylov complexity using the closed system method. First, we numerically simulate the $r_{k}$ of the two-mode squeezed state, as depicted in Fig. 3. The simulation shows oscillations during inflation, which are enhanced in RD and MD, maintaining some constant values. Notably, the different potentials produce almost identical trends for $r_{k}$. Based on the numerical results for $r_{k}$, we also simulate the corresponding Krylov complexity and Krylov entropy, as illustrated in Figs. 4 and 5. The trends of both Krylov complexity and Krylov entropy are similar across the corresponding epochs. During inflation, we observe exponential growth, consistent with our previous study Li:2024kfm . In the RD and MD phases, the trends show initial enhancement before saturating to constant values. Additionally, various inflationary models exhibit nearly the same trends for Krylov complexity and Krylov entropy, confirming the approximation that the effective mass of the inflaton is approximately constant, as indicated in Li:2024ljz .
$(c)$. Within the method of an open system, we rigorously derive the wave function of an open system using the second kind of Meixner polynomials, which we name it as the open two-mode squeezed state. This derivation is represented in Eq. (73). Throughout this process, we discover that the Lanczos coefficient $b_{n}$ and the dissipative coefficient are precisely determined by the specific Hamiltonian, leading to the conclusion that $b_{n}$ is identical across both methods. Our findings indicate that $b_{n}=\alpha n$, suggesting that our universe behaves as an entirely chaotic system. Consequently, the Lyapunov index is equivalent to $b_{n}$ when $n=2$. Figure 6 illustrates the chaos present across various epochs. During the inflationary period, chaos increases exponentially before diminishing to a certain level, only to rise again due to the preheating process. Additionally, we examine the evolution of the dissipative coefficient in RD and MD. This analysis builds on previous work in Ref. Li:2024kfm , which discusses the strong dissipative nature of the system during inflation. From Fig. 7, we can confirm that the RD and MD systems are weakly dissipative, characterized by $\mu_{2}\ll 1$.
Additionally, we derive the evolution of $r_{k}$ and $\phi_{k}$, noting that the solution for $r_{k}$ is dependent on the solution for $\phi_{k}$ due to their coupling, as described by Eq. (74). Using the wave function in Eq. (73), we calculate the Krylov complexity and Krylov entropy, which are represented in Eqs. (76) and (77). The leading order of these quantities aligns well with the formulas derived from the closed system method. In accordance with standard procedures, we need the numerical values of $r_{k}$ within the open system framework, as illustrated in Fig. 8. Unlike the closed system, there is no oscillation observed for $r_{k}$ during inflation, and the trends of Krylov complexity and Krylov entropy are depicted in Figs. 9 and 10. Compared to the closed system method, we observe that the values of Krylov complexity and Krylov entropy are significantly smaller. Recalling the definition of Krylov complexity, which pertains to the operator growth in a dynamical system, we find that a smaller operator growth suggests a decoherence-like behavior within the Krylov complexity. When dissipative effects are considered, we observe decoherence-like behavior, whether the system is weakly dissipative or not.
While we have achieved a comprehensive understanding of Krylov complexity in the early universe, many ideas still remain to be explored. Here, we outline several future directions.
Firstly, we investigate the Krylov complexity in the context of single-field inflation, and it is quite natural to extend our work into the multi-field framework Liu:2020zzv ; Liu:2019xhn ; Liu:2021rgq ; Zhang:2022bde as well as $f(R)$ gravity Liu:2018hno ; Liu:2020zlr ; Liu:2018htf . This extension allows us to explore the geometrical effects of fields and higher orders of Ricci curvature on Krylov complexity. Our Hamiltonian, given by (33), is Hermitian and includes the component of an open system. The traditional approach to handling open systems in quantum mechanics involves constructing what is known as a Lindbladian Lindblad:1975ef ; Gorini:1975nb , where the corresponding Hamiltonian is non-Hermitian. In our case, a natural method is to expand the action to higher orders with respect to the Mukhanov-Sasaki variable.
Secondly, our results indicate that the Krylov complexity in a weak dissipative system (RD or MD) will not follow the general discussions mentioned in Sec. 6.2. One possible explanation for this discrepancy is the neglect of decoherence in curvature perturbation during horizon exits. Ref. Bhattacharjee:2022lzy suggests that decoherence in circuit complexity may lead to this phenomenon of saturation. A similar concept might apply to Krylov complexity, particularly concerning the role of decoherence in the construction of the wave function.
Thirdly, we have already obtained the open two-mode squeezed state as described in (73). In light this wave function, we can explore several additional avenues. For instance, we can calculate the two-point correlation functions to derive the power spectrum. Since the derivation of this wave function is model-independent, we have the flexibility to utilize different models of this wave function to test its validity. Conversely, we can also constrain the parameters $\mu_{1}$ and $\mu_{2}$ based on data from the Cosmic Microwave Background (CMB). Furthermore, it is important to highlight that Ref. Adhikari:2022oxr indicates the potential application of the CV conjecture to Krylov complexity during inflation; however, their approach is tailored for a closed system. In contrast, our universe is fundamentally an open system, which provides us with the opportunity to further investigate the validity of the CV conjecture for Krylov complexity within the context of open systems.
Acknowledgements.
LH and KH are funded by NSFC grant NO. 12165009, Hunan Natural Science Foundation NO. 2023JJ30487 and NO. 2022JJ40340.
Appendix A The calculation of $r_{k}$ and $\phi_{k}$ within the method of open system
In this appendix, we will give the detailed calculations for $r_{k}$ and $\phi_{k}$ based on the open two-mode squeezed state,
$|\psi\rangle=\frac{\rm{sech}\ r_{k}}{1+\mu_{2}\tanh r_{k}}\sum_{n=0}^{\infty}(-1)^{n}|1-\mu_{1}^{2}|^{\frac{n}{2}}e^{2in\phi_{k}}\Bigl{(}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr{)}^{n}|n,n\rangle_{\vec{k},-\vec{k}}$ (78)
We will write down our Hamiltonian (33) as follows,
$\hat{H}=\bigl{(}k+\frac{a^{2}V_{,\phi\phi}}{2k}\bigr{)}(\hat{c}_{\vec{k}}\hat{c}_{\vec{k}}^{\dagger}+\hat{c}_{-\vec{k}}^{\dagger}\hat{c}_{-\vec{k}})+\bigl{(}\frac{a^{2}V_{,\phi\phi}}{2k}+i\frac{a^{\prime}}{a}\bigr{)}\hat{c}_{-\vec{k}}^{\dagger}\hat{c}_{\vec{k}}^{\dagger}+\bigl{(}\frac{a^{2}V_{,\phi\phi}}{2k}-i\frac{a^{\prime}}{a}\bigr{)}\hat{c}_{-\vec{k}}\hat{c}_{\vec{k}}$ (79)
[…]
$\begin{split}\hat{H}|\psi\rangle=&\ \frac{\rm sech~{}r_{k}}{1+\mu_{2}\tanh r_{k}}\sum_{n=0}^{\infty}(-1)^{n}|1-\mu_{1}^{2}|^{\frac{n}{2}}e^{2in\phi_{k}}\Bigl{(}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr{)}^{n}\Bigl{[}\bigl{(}k+\frac{a^{2}V_{,\phi\phi}}{2k}\bigr{)}(\hat{c}_{\vec{k}}\hat{c}_{\vec{k}}^{\dagger}\\ &\ +\hat{c}_{-\vec{k}}^{\dagger}\hat{c}_{-\vec{k}})|n,n\rangle_{\vec{k},-\vec{k}}+\bigl{(}\frac{a^{2}V_{,\phi\phi}}{2k}+i\frac{a^{\prime}}{a}\bigr{)}\hat{c}_{-\vec{k}}^{\dagger}\hat{c}_{\vec{k}}^{\dagger}|n,n\rangle_{\vec{k},-\vec{k}}+\bigl{(}\frac{a^{2}V_{,\phi\phi}}{2k}-i\frac{a^{\prime}}{a}\bigr{)}\hat{c}_{-\vec{k}}\hat{c}_{\vec{k}}|n,n\rangle_{\vec{k},-\vec{k}}\Bigr{]}\\ =&\ \frac{\rm sech\ r_{k}}{1+\mu_{2}\tanh r_{k}}\sum_{n=0}^{\infty}(-1)^{n}|1-\mu_{1}^{2}|^{\frac{n}{2}}e^{2in\phi_{k}}\Bigl{(}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr{)}^{n}\Bigl{[}\bigl{(}k+\frac{a^{2}V_{,\phi\phi}}{2k}\bigr{)}(2n+1)\\ &\ |n,n\rangle_{\vec{k},-\vec{k}}+\bigl{(}\frac{a^{2}V_{,\phi\phi}}{2k}+i\frac{a^{\prime}}{a}\bigr{)}(n+1)|n+1,n+1\rangle_{\vec{k},-\vec{k}}+\bigl{(}\frac{a^{2}V_{,\phi\phi}}{2k}-i\frac{a^{\prime}}{a}\bigr{)}n|n-1,n-1\rangle_{\vec{k},-\vec{k}}\Bigr{]}\\ =&\ \frac{\rm sech\ r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigl{[}\bigl{(}k+\frac{a^{2}V_{,\phi\phi}}{2k}\bigr{)}-|1-\mu_{1}^{2}|^{\frac{1}{2}}e^{2i\phi_{k}}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\bigl{(}\frac{a^{2}V_{,\phi\phi}}{2k}-i\frac{a^{\prime}}{a}\bigr{)}\Bigr{]}|0,0\rangle_{\vec{k},-\vec{k}}\\ &\ +\frac{\rm sech\ r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigl{[}\bigl{(}k+\frac{a^{2}V_{,\phi\phi}}{2k}\bigr{)}-|1-\mu_{1}^{2}|^{\frac{1}{2}}e^{2i\phi_{k}}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\bigl{(}\frac{a^{2}V_{,\phi\phi}}{2k}-i\frac{a^{\prime}}{a}\bigr{)}\Bigr{]}\\ &\ \sum_{n=1}^{\infty}(-1)^{n}|1-\mu_{1}^{2}|^{\frac{n}{2}}e^{2in\phi_{k}}\Bigl{(}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr{)}^{n}|n,n\rangle_{\vec{k},-\vec{k}}+\frac{\rm sech\ r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigl{[}\bigl{(}2k+\frac{a^{2}V_{,\phi\phi}}{k}\bigr{)}\\ &\ -|1-\mu_{1}^{2}|^{\frac{1}{2}}e^{2i\phi_{k}}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\bigl{(}\frac{a^{2}V_{,\phi\phi}}{2k}-i\frac{a^{\prime}}{a}\bigr{)}-|1-\mu_{1}^{2}|^{-\frac{1}{2}}e^{-2i\phi_{k}}\Bigl{(}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr{)}^{-1}\\ &\ \bigl{(}\frac{a^{2}V_{,\phi\phi}}{2k}+i\frac{a^{\prime}}{a}\bigr{)}\Bigr{]}\sum_{n=1}^{\infty}(-1)^{n}n|1-\mu_{1}^{2}|^{\frac{n}{2}}e^{2in\phi_{k}}\Bigl{(}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr{)}^{n}|n,n\rangle_{\vec{k},-\vec{k}}\end{split}$ (80)
Then, the right hand side is following as,
$\begin{split}i\frac{d}{d\eta}|\psi\rangle=&\ i\frac{d}{d\eta}\Bigl{[}\frac{\rm sech\ r_{k}}{1+\mu_{2}\tanh r_{k}}\sum_{n=0}^{\infty}(-1)^{n}|1-\mu_{1}^{2}|^{\frac{n}{2}}e^{2in\phi_{k}}\Bigl{(}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr{)}^{n}|n,n\rangle_{\vec{k},-\vec{k}}\Bigr{]}\\ =&\ i\Bigl{[}-\frac{(\sinh 2r_{k}+2\mu_{2}\cosh^{2}r_{k})r_{k}^{\prime}+\sinh 2r_{k}\mu_{2}^{\prime}}{2\cosh^{3}r_{k}(1+\mu_{2}\tanh r_{k})^{2}}\sum_{n=0}^{\infty}(-1)^{n}|1-\mu_{1}^{2}|^{\frac{n}{2}}e^{2in\phi_{k}}\Bigl{(}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr{)}^{n}\\ &\ |n,n\rangle_{\vec{k},-\vec{k}}\Bigr{]}+i\Bigl{[}\frac{\rm sech\ r_{k}}{1+\mu_{2}\tanh r_{k}}\sum_{n=0}^{\infty}(-1)^{n}|1-\mu_{1}^{2}|^{\frac{n}{2}}e^{2in\phi_{k}}\Bigl{(}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr{)}^{n}\\ &\ \bigl{[}2i\phi_{k}^{\prime}+\frac{2r_{k}^{\prime}}{\sin 2r_{k}(1+\mu_{2}\tanh r_{k})}+\frac{|1-\mu_{1}^{2}|^{\prime}}{2|1-\mu_{1}^{2}|}-\frac{\tanh r_{k}\mu_{2}^{\prime}}{1-\mu_{2}\tanh r_{k}}\bigr{]}|n,n\rangle_{\vec{k},-\vec{k}}\Bigr{]}\\ =&\ -i\frac{(\sinh 2r_{k}+2\mu_{2}\cosh^{2}r_{k})r_{k}^{\prime}+\sinh 2r_{k}\mu_{2}^{\prime}}{2\cosh^{3}r_{k}(1+\mu_{2}\tanh r_{k})^{2}}|0,0\rangle_{\vec{k},-\vec{k}}\\ &\ -i\frac{(\sinh 2r_{k}+2\mu_{2}\cosh^{2}r_{k})r_{k}^{\prime}+\sinh 2r_{k}\mu_{2}^{\prime}}{2\cosh^{3}r_{k}(1+\mu_{2}\tanh r_{k})^{2}}\sum_{n=1}^{\infty}(-1)^{n}|1-\mu_{1}^{2}|^{\frac{n}{2}}e^{2in\phi_{k}}\Bigl{(}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr{)}^{n}\\ &\ |n,n\rangle_{\vec{k},-\vec{k}}+\frac{\rm sech\ r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigl{(}-2\phi_{k}^{\prime}+i\frac{2r_{k}^{\prime}}{\sin 2r_{k}(1+\mu_{2}\tanh r_{k})}+i\frac{|1-\mu_{1}^{2}|^{\prime}}{2|1-\mu_{1}^{2}|}\\ &\ -i\frac{\tanh r_{k}\mu_{2}^{\prime}}{1-\mu_{2}\tanh r_{k}}\Bigr{)}\sum_{n=1}^{\infty}(-1)^{n}n|1-\mu_{1}^{2}|^{\frac{n}{2}}e^{2in\phi_{k}}\Bigl{(}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr{)}^{n}|n,n\rangle_{\vec{k},-\vec{k}}\end{split}$ (81)
[…]
$\begin{split}&\ A_{1}|0,0\rangle_{\vec{k},-\vec{k}}+A_{1}\sum_{n=1}^{\infty}(-1)^{n}|1-\mu_{1}^{2}|^{\frac{n}{2}}e^{2in\phi_{k}}\Bigl{(}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr{)}^{n}|n,n\rangle_{\vec{k},-\vec{k}}\\ &\ +A_{2}\sum_{n=1}^{\infty}(-1)^{n}n|1-\mu_{1}^{2}|^{\frac{n}{2}}e^{2in\phi_{k}}\Bigl{(}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr{)}^{n}|n,n\rangle_{\vec{k},-\vec{k}}\\ =&\ B_{1}|0,0\rangle_{\vec{k},-\vec{k}}+B_{1}\sum_{n=1}^{\infty}(-1)^{n}|1-\mu_{1}^{2}|^{\frac{n}{2}}e^{2in\phi_{k}}\Bigl{(}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr{)}^{n}|n,n\rangle_{\vec{k},-\vec{k}}\\ &\ +B_{2}\sum_{n=1}^{\infty}(-1)^{n}n|1-\mu_{1}^{2}|^{\frac{n}{2}}e^{2in\phi_{k}}\Bigl{(}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr{)}^{n}|n,n\rangle_{\vec{k},-\vec{k}}.\end{split}$ (82)
Thus, we obtain $A_{1}=B_{1}$ as
$\begin{split}&\ -i\frac{(\sinh 2r_{k}+2\mu_{2}\cosh^{2}r_{k})r_{k}^{\prime}+\sinh 2r_{k}\mu_{2}^{\prime}}{2\cosh^{3}r_{k}(1+\mu_{2}\tanh r_{k})^{2}}\\ =\frac{\rm sech\ r_{k}}{1+\mu_{2}\tanh r_{k}}&\ \Bigl{[}\bigl{(}k+\frac{a^{2}V_{,\phi\phi}}{2k}\bigr{)}-|1-\mu_{1}^{2}|^{\frac{1}{2}}e^{2i\phi_{k}}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\bigl{(}\frac{a^{2}V_{,\phi\phi}}{2k}-i\frac{a^{\prime}}{a}\bigr{)}\Bigr{]},\end{split}$ (83)
and we’re concerned about the imaginary component,
$\begin{split}&\ -i\frac{(\sinh 2r_{k}+2\mu_{2}\cosh^{2}r_{k})r_{k}^{\prime}+\sinh 2r_{k}\mu_{2}^{\prime}}{2\cosh^{3}r_{k}(1+\mu_{2}\tanh r_{k})^{2}}\\ =-\frac{\rm sech\ r_{k}}{1+\mu_{2}\tanh r_{k}}&\ \Bigl{[}|1-\mu_{1}^{2}|^{\frac{1}{2}}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\bigl{(}\frac{a^{2}V_{,\phi\phi}}{2k}\sin 2\phi_{k}-\frac{a^{\prime}}{a}\cos 2\phi_{k}\bigr{)}\Bigr{]}.\end{split}$ (84)
and therefore, we could obtain the evolution of $r_{k}$ as,
$r_{k}^{\prime}=\frac{|1-\mu_{1}^{2}|^{\frac{1}{2}}\sinh 2r_{k}\bigl{(}\frac{a^{2}V_{,\phi\phi}}{2k}\sin 2\phi_{k}-\frac{a^{\prime}}{a}\cos 2\phi_{k}\bigr{)}-\sinh 2r_{k}\mu_{2}^{\prime}}{(\sinh 2r_{k}+2\mu_{2}\cosh^{2}r_{k})}.$ (85)
The other side, we have $A_{2}=B_{2}$ showing as
$\begin{split}&\ \frac{\rm sech\ r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigl{[}-2\phi_{k}^{\prime}+i\frac{2r_{k}^{\prime}}{\sin 2r_{k}(1+\mu_{2}\tanh r_{k})}+i\frac{|1-\mu_{1}^{2}|^{\prime}}{2|1-\mu_{1}^{2}|}-i\frac{\tanh r_{k}\mu_{2}^{\prime}}{1-\mu_{2}\tanh r_{k}}\Bigr{]}\\ =&\ \frac{\rm sech\ r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigl{[}\bigl{(}2k+\frac{a^{2}V_{,\phi\phi}}{k}\bigr{)}-|1-\mu_{1}^{2}|^{\frac{1}{2}}e^{2i\phi_{k}}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\bigl{(}\frac{a^{2}V_{,\phi\phi}}{2k}-i\frac{a^{\prime}}{a}\bigr{)}\\ &\ -|1-\mu_{1}^{2}|^{-\frac{1}{2}}e^{-2i\phi_{k}}\Bigl{(}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr{)}^{-1}\bigl{(}\frac{a^{2}V_{,\phi\phi}}{2k}+i\frac{a^{\prime}}{a}\bigr{)}\Bigr{]},\end{split}$ (86)
the real component can read as,
$\begin{split}-\frac{2\phi_{k}^{\prime}\rm sech\ r_{k}}{1+\mu_{2}\tanh r_{k}}&\ =\frac{\rm sech\ r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigl{[}\bigl{(}2k+\frac{a^{2}V_{,\phi\phi}}{k}\bigr{)}\\ -\bigl{(}\frac{a^{2}V_{,\phi\phi}}{2k}\cos 2\phi_{k}+\frac{a^{\prime}}{a}\sin 2\phi_{k}\bigr{)}&\ \Bigl{(}|1-\mu_{1}^{2}|^{\frac{1}{2}}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}+|1-\mu_{1}^{2}|^{-\frac{1}{2}}(\tanh^{-1}r_{k}+\mu_{2})\Bigr{)}\Bigr{]}.\end{split}$ (87)
Finally, we could obtain the evolution equation of $\phi_{k}$ as follows,
$\begin{split}\phi_{k}^{\prime}=&\ -\bigl{(}k+\frac{a^{2}V_{,\phi\phi}}{2k}\bigr{)}+\frac{1}{2}\bigl{(}\frac{a^{2}V_{,\phi\phi}}{2k}\cos 2\phi_{k}+\frac{a^{\prime}}{a}\sin 2\phi_{k}\bigr{)}\\ &\ \Bigl{[}|1-\mu_{1}^{2}|^{\frac{1}{2}}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}+|1-\mu_{1}^{2}|^{-\frac{1}{2}}(\tanh^{-1}r_{k}+\mu_{2})\Bigr{]}.\end{split}$ (88)
One can easily check that Eqs. (85) and (88) could nicely recover into Eq. (35) as $\mu_{1}=\mu_{2}=0$.